OpenStudy (anonymous):

solve: sqrt5x+21=x+3 solution is x?

OpenStudy (anonymous):

\[\sqrt{5x+21}=x+3\]

OpenStudy (amistre64):

we under the sqrt by ^2 both sides: 5x +21 = x^2 +6x +9 ; -5x and -21 to get 0 = x^2 +x -12 ; factor the quad to get: x = -4,3

OpenStudy (amistre64):

dbl chk because this way can produce false results

OpenStudy (amistre64):

x = 3 works out right,the -4 is just a by product of the method

OpenStudy (anonymous):

yeah it is only asking for one answer

OpenStudy (amistre64):

then id go with 3 :) sqrt(5(3) +21) = 3+3 sqrt(36) = 6

OpenStudy (anonymous):

3

OpenStudy (anonymous):

3 is the answer coz your solution must satisfy the given equation

OpenStudy (anonymous):

well we can write this equetion like that 5x+21 = x ^2 + 6x + 9 x + x ^ 2 = 12 x=3 ^^